White dot for spacing only
The Dice Project


Operational Meeting

Minutes of the meeting held on Wednesday 28th April 2010
Room 4.31, Informatics Forum

Present:
Ross Armstrong, Richard Bell, Toby Blake, Neil Brown, Roger Burroughes, Chris Cooke, Graham Dutton, Stephen Quinney, Iain Rae, Gordon Reid, George Ross, Alastair Scobie, Craig Strachan (convener and minutes)
  1. Apologies for absence

    Alison, Lindsey, Tim, Ian

  2. Minutes of the last meeting.

    These were accepted.

  3. Actions.

    Actions completed:

    Actions deferred:

    Actions added:

  4. Report from Computing Executive Group

  5. Reports from units.

  6. Topics for discussion

    1. School's new Health and Safety web site doesn't appear to cover our activities

      It's clear from a study of the new Health and Safety web site that it doesn't cover a large number of activities which the computing staff are required to carry out. Amongst the issues identified were

      • Working with floor tiles
      • The presence of three phase in server rooms
      • Lifting heavy hardware
      • handling UPSes, both in terms of their weight and the power they contain
      • Working in comms closets and other confined spaces
      • Noise levels in server rooms
      • working alone in server rooms

      In addition, it was pointed out that none of the risk assessment forms include a computing staff category.

      It was agreed that Alison, our representative on the Health and Safety committee should compile a list of our concerns and present it at the next Health and Safety committee meeting.

    2. Usage of KB server room

      George pointed out that when the infrastructure had been set up for the KB server room, it had been designed on the assumption that the kit at KB would only be used for disaster recovery. At the moment, there is some commodity date, mainly AFS RW volumes stored on KB servers. Craig stated that the intention was to move all AFS RW volumes downtown in the near future. This does mean that we will see considerable network traffic between the central area and KB every night as the RO mirror volumes are updated and we will need to monitor how the KB infrastructure copes with this. We have a deadline of August for deciding whether we need to upgrade the network connection to KB. It was pointed out that Physics will shortly be installing DR kit in the Forum server room.

  7. AOCB

    1. ssh server update

      Stephen reported that we are now running fail2ban on the ssh servers. When login failures from a given IP address exceed a certain limit, further connections from that address are blocked via the hosts.deny file for a given period of time. Stephen is monitoring the behavior of this script to make sure that the various time limits have a reasonable value. It was agreed that user support should send a suitably vague message to the user base announcing that this measure had been put in place.

  8. Next meeting

    The next meeting will be on Wednesday 12th May in room 4.31.


 : Operational : Meetings 

Mini Informatics Logo - Link to Main Informatics Page
Please contact us with any comments or corrections.
Unless explicitly stated otherwise, all material is copyright The University of Edinburgh
Spacing Line